OpenStudy (anonymous):

sqrt(3x)+3sqrt(3x)-2sqrt(3x)-10

OpenStudy (mertsj):

You have three terms with similar radicals. Add their coefficients.

OpenStudy (anonymous):

I keep getting 2(sqrt(3x)-5)

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

\[\sqrt{3x}+3\sqrt{3x}-2\sqrt{3x}-10=\sqrt{3x}(1+3-2)-10=\sqrt{3x}(2)-10=2\sqrt{3x}-10\]

OpenStudy (mertsj):

-10 cut off at end

OpenStudy (anonymous):

So I leave the -10 or keep it?

OpenStudy (mertsj):

keep it

OpenStudy (anonymous):

Ok thank you , your awesome

OpenStudy (mertsj):

yw